OpenStudy (anonymous):

Medal A 250 mL beaker can hold 0.192 kg of acetone. What is the density of this substance, in g/mL? Round your answer to 3 significant digits

OpenStudy (jhannybean):

Even though this should be in the chemistry section..

OpenStudy (jhannybean):

\[\sf 0.192~kg~C_3H_6O ~\times~\frac{10^3~g~C_3H_6O}{1~kg~C_3H_6O} =~?~g~C_3H_6O\]\[\sf =\frac{g~C_3H_6O}{250~\times~10^{-3}~L}\]

OpenStudy (anonymous):

how would you solve this

OpenStudy (jhannybean):

Follow my steps